OpenStudy (anonymous):

A flagpole casts a 15 ft shadow at the same time Jon casts a 5 ft shadow. If Jon is 6 ft tall, how tall is the flagpole?

OpenStudy (anonymous):

 6/5 = x/15 ; i.e. x=18 feet